1) Given 2 jobs: Job A which pays $10,000 in the 1st year and $57,000 in the 2nd year Job B which requires attending a seminar in the first year costing $13,000 (to be paid out of pocket) then making $88,000 in the 2nd year. If a person has a discount rate of 9% which one would he choose?

2)Given 2 jobs: Job A which pays $10,000 in the 1st year and $57,000 in the 2nd year Job B which requires attending a seminar in the first year costing $13,000 (to be paid out of pocket) then making $88,000 in the 2nd year. If a person has a discount rate of 42% which one would he choose?
Answer #1

Q1.

Job A; Earnings = $10000 (Year 1) + PV ($57000) = 10000 + 57000/1.09 = $62293 { PV = FV/1+r }

Job B: Earnings = -$13000 (Year 1) + PV ($88000 ) = -13000 + 88000/1.09 = $70018

Hence Job B is more profitable

Q2.

Job A; Earnings = $10000 (Year 1) + PV ($57000) = 10000 + 57000/1.42 = $52140 { PV = FV/1+r }

Job B: Earnings = -$13000 (Year 1) + PV ($88000 ) = -13000 + 88000/1.42 = $48971

Hence Job A is more profitable